# From Ferminet to PINN. Connections between neural network-based algorithms for high-dimensional Schrödinger Hamiltonian

### v1

@@ 3.1 Neural network optimization
    $\left.\begin{array}[]{lcl}\nabla_{\vartheta}\mathcal{M}^{(\textrm{PINN})}_{i}(\vartheta)&=&2\big{\langle}\nabla_{\vartheta}\widehat{\psi}(\vartheta;\cdot),\widehat{\psi}(\vartheta;\cdot)-(1-dtH)\widehat{\phi}_{i-1}\big{\rangle}\\ &=&2\big{\langle}\nabla_{\vartheta}\widehat{\psi}(\vartheta;\cdot),\widehat{\psi}(\vartheta;\cdot)-\widehat{\phi}_{i-1}\big{\rangle}-2dt\big{\langle}\nabla_{\vartheta}\widehat{\psi}(\vartheta;\cdot),H\widehat{\phi}_{i-1}\big{\rangle}\\ &=&-2\big{\langle}\nabla_{\vartheta}\widehat{\psi}(\vartheta;\cdot),\widehat{\psi}(\vartheta;\cdot)-\widehat{\phi}_{i-1}-dtH\widehat{\phi}_{i-1}\big{\rangle}+O(dt^{2})\,.\end{array}\right.$
@@
    $\vartheta_{k+1}=\vartheta_{k}-\cfrac{2\nu_{k}}{M}\sum_{\alpha}\big{\{}\nabla_{\vartheta}\widehat{\psi}(\vartheta_{k};R_{i}^{\alpha})\big{[}\widehat{\psi}(\vartheta;R_{i}^{\alpha})-\widehat{\phi}_{i-1}(R_{i}^{\alpha})-dtH\widehat{\phi}_{i-1}(R_{i}^{\alpha})\big{]}\big{\}}+O(dt^{2})\,.$ (64)
@@
    $\left.\begin{array}[]{lcl}\partial_{t}f(R,t)&=&K(f_{i-1})(R),\,\,\,(R,t)\in Q_{i},\\ f(R,t_{i-1})&=&f_{i-1}(R),\,\,\,R\in\mathbb{R}^{3N}\,.\end{array}\right.$ (67)
@@
    $\left.\begin{array}[]{lcl}\mathcal{L}^{(\textrm{PINN})}_{i}(\theta)&=&\big{\|}\widehat{f}(\theta;\cdot)-\widehat{f}_{i-1}-\cfrac{dt}{2}\nabla_{R}^{2}\widehat{f}_{i-1}+dt\nabla_{R}(\widehat{V}_{i-1}\widehat{f}_{i-1})-dt\widehat{S}_{i-1}\widehat{f}_{i-1}\big{\|}^{2}_{L^{2}(\mathbb{R}^{3N})}\\ &=&\big{\|}\widehat{f}(\theta;\cdot)-\widehat{f}_{i-1}-dtK(\widehat{f}_{i-1})\big{\|}^{2}_{L^{2}(\mathbb{R}^{3N})}\end{array}\right.$
@@
    $\left.\begin{array}[]{lcl}\nabla_{\theta}\mathcal{L}^{(\textrm{PINN})}_{i}(\theta)&=&2\big{\langle}\nabla_{\theta}\widehat{f}(\theta;\cdot),\widehat{f}(\theta;\cdot)-\widehat{f}_{i-1}-dtK(\widehat{f}_{i-1})\big{\rangle}\,.\end{array}\right.$
@@
    $\widehat{f}(\theta;\cdot)=\widehat{\phi}_{i-1}\widehat{\phi}(\theta;\cdot),\,\,\,\widehat{f}_{i-1}=\widehat{\phi}_{i-1}^{2}\,.$
    Then, from (49)
    $\left.\begin{array}[]{lcl}\nabla_{\theta}\mathcal{L}^{(\textrm{PINN})}_{i}(\theta)&=&2\big{\langle}\widehat{\phi}_{i-1}\nabla_{\theta}\widehat{\phi}(\theta;\cdot),\widehat{\phi}_{i-1}\widehat{\phi}(\theta;\cdot)-\widehat{\phi}^{2}_{i-1}-dtK(\widehat{\phi}^{2}_{i-1})\big{\rangle}\\ &=&2\big{\langle}\widehat{\phi}_{i-1}\nabla_{\theta}\widehat{\phi}(\theta;\cdot),\widehat{\phi}_{i-1}\widehat{\phi}(\theta;\cdot)-\widehat{\phi}^{2}_{i-1}-\cfrac{dt}{2}\nabla_{R}^{2}\widehat{\phi}^{2}_{i-1}+dt\nabla_{R}(\widehat{V}_{i-1}\widehat{\phi}^{2}_{i-1})\\ &&-dt\widehat{S}_{i-1}\widehat{\phi}^{2}_{i-1}\big{\rangle}\\ &=&2\big{\langle}\widehat{\phi}_{i-1}\nabla_{\theta}\widehat{\phi}(\theta;\cdot),\widehat{\phi}_{i-1}\widehat{\phi}(\theta;\cdot)-\widehat{\phi}^{2}_{i-1}-dt|\nabla_{R}\widehat{\phi}_{i-1}|^{2}-dt\widehat{\phi}_{i-1}\nabla^{2}_{R}\widehat{\phi}_{i-1}\\ &&+dt\widehat{\phi}^{2}_{i-1}\nabla_{R}\widehat{V}_{i-1}+2dt\widehat{\phi}_{i-1}\nabla_{R}\widehat{\phi}_{i-1}-dt\widehat{S}_{i-1}\widehat{\phi}^{2}_{i-1}\big{\rangle}\\ &=&2\big{\langle}\widehat{\phi}^{2}_{i-1}\nabla_{\theta}\widehat{\phi}(\theta;\cdot),\widehat{\phi}(\theta;\cdot)-\widehat{\phi}_{i-1}-dt\widehat{\phi}_{i-1}^{-1}|\nabla_{R}\widehat{\phi}_{i-1}|^{2}-dt\nabla_{R}\widehat{\phi}_{i-1}\\ &&+dt\widehat{\phi}_{i-1}\nabla_{R}\widehat{V}_{i-1}+2dt\nabla^{2}_{R}\widehat{\phi}_{i-1}-dt\widehat{S}_{i-1}\widehat{\phi}_{i-1}\big{\rangle}\\ &=&2\big{\langle}\widehat{\phi}^{2}_{i-1}\nabla_{\theta}\widehat{\phi}(\theta;\cdot),\widehat{\phi}(\theta;\cdot)-\widehat{\phi}_{i-1}-dtH\widehat{\phi}_{i-1}\big{\rangle}\,.\par\end{array}\right.$
@@
    $\left.\begin{array}[]{lcl}\nabla_{\theta}\mathcal{L}^{(\textrm{PINN})}_{i}(\theta)&=&2\big{\langle}\nabla_{\theta}\widehat{\phi}(\theta;\cdot),\widehat{\phi}(\theta;\cdot)-\widehat{\phi}_{i-1}-dtH\widehat{\phi}_{i-1}\big{\rangle}_{\widehat{\phi}_{i-1}}\,.\end{array}\right.$
@@
    $\theta_{k+1}=\theta_{k}-\cfrac{2\nu_{k}}{M}\sum_{\alpha}\big{\{}\sum_{\alpha}\widehat{\phi}^{2}_{i-1}(R_{i}^{\alpha})\nabla_{\theta}\widehat{\phi}(\theta;R_{i}^{\alpha})\big{[}\widehat{\phi}(\theta;(R_{i}^{\alpha}))-\widehat{\phi}_{i-1}(R_{i}^{\alpha})-dtH\widehat{\phi}_{i-1}(R_{i}^{\alpha})\big{]}\big{\}}\,.$ (72)
@@
    and where the corresponding gradient descent method is initialized with the solution to the DMC solution at time $t_{i-1}$ (which differs from $\widehat{\phi}_{i-1}$). The converged solution is then denoted by $\widehat{\phi}_{i-1}$. Let us consider VMC///DMC on $Q_{i}$ assuming that DMC approximates
@@
    for $\langle\widehat{\phi}_{i-1},\widehat{\phi}_{i-1}\rangle=1$. Hence
@@
    We have (for $dt$ small enough)
    $\left.\begin{array}[]{lcl}\langle\psi(\vartheta;\cdot),\psi(\vartheta;\cdot)\rangle&=&1-dt\langle\widehat{\phi}_{i-1},H(\psi(\vartheta;\cdot)+\widehat{\phi}_{i-1})\rangle+\cfrac{dt^{2}}{4}\|H\big{(}\psi(\vartheta;\cdot)+\widehat{\phi}_{i-1}\big{)}\|^{2}\\ &=&1-dt+2dt\langle H\widehat{\phi}_{i-1},\widehat{\phi}_{i-1}\rangle-dt\langle\psi(\vartheta;\cdot),H\widehat{\phi}_{i-1}\rangle\\ &&+\cfrac{dt^{2}}{4}\|H\big{(}\psi(\vartheta;\cdot)+\widehat{\phi}_{i-1}\big{)}\|^{2}\\ &=&1-dt+dt\langle H\widehat{\phi}_{i-1},\widehat{\phi}_{i-1}\rangle+O(dt^{2})\,.\end{array}\right.$
@@
    $\left.\begin{array}[]{lcl}\langle\psi(\vartheta;\cdot),\psi(\vartheta;\cdot)\rangle^{-1}&=&1+dt-dt\langle H\widehat{\phi}_{i-1},\widehat{\phi}_{i-1}\rangle+O(dt^{2})\,,\end{array}\right.$
    and
    $\left.\begin{array}[]{lcl}\cfrac{\big{\langle}H\widehat{\psi}(\vartheta;\cdot),\widehat{\psi}(\vartheta;\cdot)\big{\rangle}}{\big{\langle}\psi(\vartheta;\cdot),\psi(\vartheta;\cdot)\big{\rangle}}&=&-2\big{\langle}\widehat{\psi}(\vartheta;\cdot),H\widehat{\phi}_{i-1}\big{\rangle}-2-3\langle H\widehat{\phi}_{i-1},\widehat{\phi}_{i-1}\rangle+O(dt)\,.\end{array}\right.$
@@
    $\left.\begin{array}[]{lcl}\nabla_{\theta}\mathcal{R}_{i}(\vartheta)&=&-2\big{\langle}\nabla_{\vartheta}\widehat{\psi}(\vartheta;\cdot),H\widehat{\phi}_{i-1}\big{\rangle}+O(dt)\\ &=&\cfrac{1}{dt}\big{\langle}\nabla_{\vartheta}\widehat{\psi}(\vartheta;\cdot),\widehat{\psi}(\vartheta;\cdot)-\widehat{\phi}_{i-1}-dtH\widehat{\phi}_{i-1}\big{\rangle}+O(dt)\,.\end{array}\right.$
@@
    $\vartheta_{k+1}=\vartheta_{k}+\cfrac{2\nu_{k}}{M}\sum_{\alpha}\big{\{}\nabla_{\vartheta}\widehat{\psi}(\vartheta_{k};R_{i}^{\alpha})\big{[}\widehat{\psi}(\vartheta_{k};R_{i}^{\alpha})-\widehat{\phi}_{i-1}(R_{i}^{\alpha})-dtH\widehat{\phi}_{i-1}(R_{i}^{\alpha})+O(dt^{2})\big{]}\big{\}}\,.$ (79)
@@ 3.2 Conclusion
Recall that the “direct” N-VDMC as described in Section 1.4 (see [26]) and unlike the modified version used above, does not require to the use of VMC at each time iteration $t_{i}$ and as a consequence is much less computationally complex. However, PINN-like algorithms do not necessarily require the discretization of the time-derivative. Instead we can search for a neural network on $Q_{T}=\mathbb{R}^{3N}\times[0,T]$ (for $T$ large enough) with $f_{0}=\psi^{2}_{T}$ given, for a neural network $\widehat{f}:(\theta;R,t)\in\Theta\times Q\rightarrow\mathbb{R}$ (and designed such that $\widehat{f}:(\theta;R,0)=f_{0}(R)$) by minimizing
@@
In principle, the computation is performed once for all over $Q$, hence avoiding the use of N-VMC at intermediate times $t_{i}$. In practice, it may however be necessary to decompose the time-domain $[0,T]$ in several subdomains ($[T_{i},T_{i+1}]$ of length larger than $dt$) used in DMC to avoid the computation of a null network.
@@ 4 Conclusion
Ferminet and PINN, Deep-Ritz have become very popular libraries and algorithms which were developed independently to each other in different fields of research. In this manuscript we have established non-trivial bridges between these different methodologies. We think that these connections may be beneficial for future developments in the different involved scientific communities. Naturally, PINN algorithms for solving (2) are more computationally complex than DMC-like algorithms, as they approximate PDE solutions, unlike DMC algorithms which only evaluate “trajectories”. In future works, we plan to develop efficient PINN-like algorithms for eigenvalue problems in the spirit of DMC algorithms. Data-driven PINN algorithms will be explored for high-dimensional quantum chemistry evolution problems, such as laser-molecule interactions.

### v2

@@ 3.1 Neural network optimization
    $\left.\begin{array}[]{lcl}\nabla_{\vartheta}\mathcal{M}^{(\textrm{PINN})}_{i}(\vartheta)&=&2\big{\langle}\nabla_{\vartheta}\widehat{\psi}(\vartheta;\cdot),\widehat{\psi}(\vartheta;\cdot)-(1-dtH)\widehat{\phi}_{i-1}\big{\rangle}\,.\end{array}\right.$
@@
    $\vartheta_{k+1}=\vartheta_{k}-\cfrac{2\nu_{k}}{M}\sum_{\alpha}\big{\{}\nabla_{\vartheta}\widehat{\psi}(\vartheta_{k};R_{i}^{\alpha})\big{[}\widehat{\psi}(\vartheta;R_{i}^{\alpha})-\widehat{\phi}_{i-1}(R_{i}^{\alpha})+dtH\widehat{\phi}_{i-1}(R_{i}^{\alpha})\big{]}\big{\}}\,,$ (64)
    with learning rate $\nu_{k}$.
@@
    $\left.\begin{array}[]{lcl}\partial_{t}f(R,t)&=&-K(f_{i-1})(R),\,\,\,(R,t)\in Q_{i},\\ f(R,t_{i-1})&=&f_{i-1}(R),\,\,\,R\in\mathbb{R}^{3N}\,.\end{array}\right.$ (67)
@@
    $\left.\begin{array}[]{lcl}\mathcal{L}^{(\textrm{PINN})}_{i}(\theta)&=&\big{\|}\widehat{f}(\theta;\cdot)-\widehat{f}_{i-1}-\cfrac{dt}{2}\nabla_{R}^{2}\widehat{f}_{i-1}+dt\nabla_{R}(\widehat{V}_{i-1}\widehat{f}_{i-1})-dt\widehat{S}_{i-1}\widehat{f}_{i-1}\big{\|}^{2}_{L^{2}(\mathbb{R}^{3N})}\\ &=&\big{\|}\widehat{f}(\theta;\cdot)-\widehat{f}_{i-1}+dtK(\widehat{f}_{i-1})\big{\|}^{2}_{L^{2}(\mathbb{R}^{3N})}\end{array}\right.$
@@
    $\left.\begin{array}[]{lcl}\nabla_{\theta}\mathcal{L}^{(\textrm{PINN})}_{i}(\theta)&=&2\big{\langle}\nabla_{\theta}\widehat{f}(\theta;\cdot),\widehat{f}(\theta;\cdot)-\widehat{f}_{i-1}+dtK(\widehat{f}_{i-1})\big{\rangle}\,.\end{array}\right.$
@@
    $\widehat{f}(\theta;\cdot)=\widehat{\phi}_{i-1}\widehat{\phi}(\theta;\cdot),\,\,\,\widehat{f}_{i-1}=\widehat{\phi}_{i-1}^{2}\,,$
    then, from (49) we get
    $\left.\begin{array}[]{lcl}\nabla_{\theta}\mathcal{L}^{(\textrm{PINN})}_{i}(\theta)&=&2\big{\langle}\widehat{\phi}_{i-1}\nabla_{\theta}\widehat{\phi}(\theta;\cdot),\widehat{\phi}_{i-1}\widehat{\phi}(\theta;\cdot)-\widehat{\phi}^{2}_{i-1}+dtK(\widehat{\phi}^{2}_{i-1})\big{\rangle}\\ &=&2\big{\langle}\widehat{\phi}^{2}_{i-1}\nabla_{\theta}\widehat{\phi}(\theta;\cdot),\widehat{\phi}(\theta;\cdot)-\widehat{\phi}_{i-1}+dtH\widehat{\phi}_{i-1}\big{\rangle}\,.\par\end{array}\right.$
@@
    $\left.\begin{array}[]{lcl}\nabla_{\theta}\mathcal{L}^{(\textrm{PINN})}_{i}(\theta)&=&2\big{\langle}\nabla_{\theta}\widehat{\phi}(\theta;\cdot),\widehat{\phi}(\theta;\cdot)-\widehat{\phi}_{i-1}+dtH\widehat{\phi}_{i-1}\big{\rangle}_{\widehat{\phi}_{i-1}}\,.\end{array}\right.$
@@
    $\theta_{k+1}=\theta_{k}-\cfrac{2\nu_{k}}{M}\sum_{\alpha}\big{\{}\sum_{\alpha}\widehat{\phi}^{2}_{i-1}(R_{i}^{\alpha})\nabla_{\theta}\widehat{\phi}(\theta;R_{i}^{\alpha})\big{[}\widehat{\phi}(\theta;(R_{i}^{\alpha}))-\widehat{\phi}_{i-1}(R_{i}^{\alpha})+dtH\widehat{\phi}_{i-1}(R_{i}^{\alpha})\big{]}\big{\}}\,,$ (72)
    with learning rate $\nu_{k}$.
@@
    and where the corresponding gradient descent method is initialized with the solution to the DMC solution at time $t_{i-1}$ (which differs from $\widehat{\phi}_{i-1}$). The converged solution is then denoted by $\widehat{\phi}_{i-1}$. Let us consider VMC///DMC on $Q_{i}$ assuming that $\widehat{\psi}$ (from DMC) is such that
@@
    and $\langle\widehat{\phi}_{i-1},\widehat{\phi}_{i-1}\rangle=1$.
    We have (for $dt$ small enough)
    $\left.\begin{array}[]{lcl}\langle\psi(\vartheta;\cdot),\psi(\vartheta;\cdot)\rangle&=&1-dt\langle\widehat{\phi}_{i-1},H(\psi(\vartheta;\cdot)+\widehat{\phi}_{i-1})\rangle+\cfrac{dt^{2}}{4}\|H\big{(}\psi(\vartheta;\cdot)+\widehat{\phi}_{i-1}\big{)}\|^{2}\\ &=&1-dt\langle H\widehat{\phi}_{i-1},\widehat{\phi}_{i-1}\rangle-dt\langle\psi(\vartheta;\cdot),H\widehat{\phi}_{i-1}\rangle+O(dt^{2})\\ &=&1-2dt\langle\psi(\vartheta;\cdot),H\widehat{\phi}_{i-1}\rangle+O(dt^{2})\,.\end{array}\right.$
    Then $\nabla_{\theta}\langle\psi(\vartheta;\cdot),\psi(\vartheta;\cdot)\rangle=-dt\langle\nabla_{\theta}\psi(\vartheta;\cdot),H\widehat{\phi}_{i-1}\rangle+O(dt^{2})$. Hence
@@
    $\left.\begin{array}[]{lcl}\cfrac{\big{\langle}H\widehat{\psi}(\vartheta;\cdot),\widehat{\psi}(\vartheta;\cdot)\big{\rangle}}{\big{\langle}\psi(\vartheta;\cdot),\psi(\vartheta;\cdot)\big{\rangle}}&=&4\langle\psi(\vartheta;\cdot),H\widehat{\phi}_{i-1}\rangle-2\big{\langle}\widehat{\psi}(\vartheta;\cdot),H\widehat{\phi}_{i-1}\big{\rangle}\\ &&-\big{\langle}H\widehat{\phi}_{i-1},\widehat{\phi}_{i-1}\big{\rangle}+O(dt)\\ &=&2\big{\langle}\widehat{\psi}(\vartheta;\cdot),H\widehat{\phi}_{i-1}\big{\rangle}-\langle H\widehat{\phi}_{i-1},\widehat{\phi}_{i-1}\rangle+O(dt)\,.\end{array}\right.$
@@
    $\left.\begin{array}[]{lcl}\nabla_{\theta}\mathcal{R}_{i}(\vartheta)&=&2\big{\langle}\nabla_{\vartheta}\widehat{\psi}(\vartheta;\cdot),H\widehat{\phi}_{i-1}\big{\rangle}+O(dt)\\ &=&-\cfrac{2}{dt}\big{\langle}\nabla_{\vartheta}\widehat{\psi}(\vartheta;\cdot),\widehat{\psi}(\vartheta;\cdot)-\widehat{\phi}_{i-1}+dtH\widehat{\psi}(\vartheta;\cdot)\big{\rangle}+O(dt)\,.\end{array}\right.$
@@
    $\vartheta_{k+1}=\vartheta_{k}-\cfrac{2\nu_{k}}{M}\sum_{\alpha}\big{\{}\nabla_{\vartheta}\widehat{\psi}(\vartheta_{k};R_{i}^{\alpha})\big{[}\widehat{\psi}(\vartheta_{k};R_{i}^{\alpha})-\widehat{\phi}_{i-1}(R_{i}^{\alpha})+dtH\widehat{\phi}_{i-1}(R_{i}^{\alpha})+O(dt^{2})\big{]}\big{\}}\,,$ (78)
    with learning rate $\nu_{k}$.
@@ 3.2 Conclusion
Recall that the “direct” N-VDMC as described in Section 1.4 (see [27]) and unlike the modified version used above, does not require to the use of VMC at each time iteration $t_{i}$ and as a consequence is much less computationally complex. Notice that PINN-like algorithms do not necessarily require the discretization of the time-derivative. Instead we can search for a neural network on $Q_{T}=\mathbb{R}^{3N}\times[0,T]$ (for $T$ large enough) with $f_{0}=\psi^{2}_{T}$ given, for a neural network $\widehat{f}:(\theta;R,t)\in\Theta\times Q\rightarrow\mathbb{R}$ (and designed such that $\widehat{f}:(\theta;R,0)=f_{0}(R)$) by minimizing
@@
In principle, the computation is performed once for all over $Q$, hence avoiding the use of N-VMC at intermediate times $t_{i}$. In practice, it may still however be necessary to decompose the time-domain $[0,T]$ in several subdomains ($[T_{i},T_{i+1}]$ of length larger than $dt$) used in DMC to avoid the computation of a null network.
@@ 4 Conclusion
NQS, Ferminet and PINN, Deep-Ritz have become very popular libraries and algorithms which were developed independently to each other in different fields of research. In this manuscript we have established non-trivial bridges between these different methodologies. We think that these connections may be beneficial for future developments in the different involved scientific communities. Naturally, PINN algorithms for solving (2) are more computationally complex than DMC-like algorithms, as they approximate PDE solutions, unlike DMC algorithms which only evaluate “trajectories”. In future works, we plan to develop efficient PINN-like algorithms for eigenvalue problems in the spirit of DMC algorithms. Data-driven PINN algorithms will be explored for high-dimensional quantum chemistry evolution problems (for instance in the spirit of [21]), such as laser-molecule interactions.